# Two tales for a geometric Jensen–Shannon divergence

### v1

@@ 1 Introduction
Besides information theory, the JSD also plays an important role in machine learning [21, 11, 36]. However, one drawback that refrains its use in practice is that the JSD between two Gaussian distributions is not known in closed-form since no analytic formula is known for the differential entropy of a two-component Gaussian mixture [22], and thus the JSD needs to be numerically approximated in practice by various methods.
@@
In general, the log-mixture term integrand makes it difficult to compute in closed-form the differential entropy of mixture of densities although some exceptions may exist (e.g., the JSD between Cauchy distributions is available in closed-form [29]).
The geometric G-JSD was defined in [25] as follows:
Definition 2 (G-JSD)
$D_{\mathrm{JS}}^{G}(p_{1},p_{2}){:=}\frac{1}{2}\,\left(D_{\mathrm{KL}}(p_{1}:(p_{1}p_{2})^{G})+D_{\mathrm{KL}}(p_{2}:(p_{1}p_{2})^{G})\right),$
where $(p_{1}p_{2})^{G}(x)=\frac{\sqrt{p_{1}(x)p_{2}(x)}}{\int\sqrt{p_{1}(x)p_{2}(x)}\,{\mathrm{d}\mu}}$ is the (normalized) geometric mixture of $p_{1}$ and $p_{2}$.
To contrast with the approximation limitation of the JSD between Gaussians, one advantage of the geometric Jensen–Shannon divergence (G-JSD) is that it admits a closed-form expression between Gaussian distributions [25]. However, the G-JSD is not anymore bounded and thus requires matching support of the densities. The G-JSD formula between Gaussian distributions has been used in several scenarii. See [8, 7, 20, 18, 23, 33, 39, 34, 37, 13]) for a few use cases.
The paper is organized as follows: We first give an alternative definition of the M-JSD in §2 (Definition 3) which extends to positive measures and do not require normalization of geometric mixtures. We call this new divergence the extended M-JSD and we compare the two types of geometric JSDs when dealing with probability measures. In §4, we show that these normalized/extended $M$-JSD divergences can be interpreted as regularizations of the Jensen–Shannon divergence, and exhibit several bounds. We report related closed-form formula for the G-JSD and extended G-JSD between two Gaussian distributions in Section 3. Finally, we summarize the main results in the concluding section §5.
@@ 2 A novel definition G-JSD extended to positive measures
  • First, we replace the KLD by the extended KLD (see [30] for more general extended $f$-divergences) between positive densities $q_{1}\in M_{\mu}^{+}$ and $q_{2}\in M_{\mu}^{+}$ instead of normalized densities:
    $D_{{\mathrm{KL}}^{+}}(q_{1},q_{2}){:=}\int\left(q_{1}\log\frac{q_{1}}{q_{2}}+q_{2}-q_{1}\right)\,{\mathrm{d}\mu},$ (6)
    and,
  • Second, we consider unnormalized $M$-mixture densities ${(q_{1}q_{2})}_{\alpha}^{\tilde{M}}(x){:=}{M_{\alpha}(q_{1}(x),q_{2}(x))}$ where we use the $\tilde{M}$ notation to indicate that the $M$-mixture is not normalized, instead of normalized densities ${(q_{1}q_{2})}_{\alpha}^{M}(x)$.
The extended KLD can be interpreted as a pointwise integral of a scalar Bregman divergence obtained for the negative Shannon entropy generator [16]. This proves that $D_{{\mathrm{KL}}^{+}}(q_{1},q_{2})\geq 0$ with equality if and only if $q_{1}=q_{2}$ $\mu$-almost everywhere. Notice that $D_{\mathrm{KL}}(q_{1},q_{2})$ may be negative when $q_{1}$ and/or $q_{2}$ are not normalized to probability densities but we always have $D_{{\mathrm{KL}}^{+}}(q_{1},q_{2})\geq 0$.
@@
As a side remark, it is preferable in practice to estimate the KLD between $p_{1}$ and $p_{2}$ by Monte Carlo methods using Eq. 6 instead of Eq. 1 in order to guarantee the non-negativeness of the KLD (Gibbs’ inequality). Indeed, the sampling of $s$ samples $x_{1},\ldots,x_{s}$, defines two unnormalized distributions $q_{1}(x)=\frac{1}{s}\sum_{i=1}^{s}p_{1}(x)\delta_{x_{i}}(x)$ and $q_{2}(x)=\frac{1}{s}\sum_{i=1}^{s}p_{2}(x)\delta_{x_{i}}(x)$ where $\delta_{x_{i}}(x)=\left\{\begin{array}[]{ll}1,&\mbox{if $x=x_{i}$}\cr 0,&\mbox{otherwise}\end{array}\right$..
@@
where $Z(q){:=}\int q\,{\mathrm{d}\mu}$ is the normalization factor of the positive density $q$. The divergence $\tilde{D}$ is said projective because we have for all $\lambda_{1}>0,\lambda_{2}>0$, the property that $\tilde{D}(\lambda_{1}q_{1},\lambda_{2}q_{2})=\tilde{D}(q_{1},q_{2})=D^{+}(p_{1},p_{2})$ where $p_{i}=\frac{q_{i}}{Z(q_{i})}$ are the normalized densities. The projective Kullback–Leibler divergence $\tilde{D}_{\mathrm{KL}}$ is thus another projective extension of the KLD to non-normalized densities which coincide with the KLD for probability densities. But the projective KLD is different from the extended KLD of Eq. 6 and furthermore we have $\tilde{D}_{\mathrm{KL}}(q_{1},q_{2})=0$ if and only if $q_{1}=\lambda q_{2}$ $\mu$-almost everywhere for some $\lambda>0$.
@@
When $\beta=\frac{1}{2}$, we simplify the notation $\mathrm{JS}_{D^{+}}^{\tilde{M}_{\alpha}}$ for short, and furthermore when $\alpha=\frac{1}{2}$, we write $\mathrm{JS}_{D^{+}}^{\tilde{M}}$ to avoid burden notations. Let $D_{\mathrm{JS}^{+}}^{M}=\mathrm{JS}_{\mathrm{JS}^{+}}^{\tilde{M}_{\frac{1}{2}},\frac{1}{2}}$.
@@
$D_{{\mathrm{KL}}^{+}}(q_{1},q_{2}){:=}D_{\mathrm{KL}}(q_{1},q_{2})+\int\left(q_{2}-q_{1}\right)\,{\mathrm{d}\mu},$
although $q_{1}$ and $q_{2}$ may not need to be normalized in the $D_{\mathrm{KL}}$ term (which can then be potentially negative). Letting $Z(q_{i}){:=}\int q_{i}\,{\mathrm{d}\mu}$ be the total mass of $q_{i}$, we have
$D_{{\mathrm{KL}}^{+}}(q_{1},q_{2})=D_{\mathrm{KL}}(q_{1},q_{2})+Z(q_{2})-Z(q_{1}).$ (8)
Let ${\tilde{m}}_{\alpha}=M_{\alpha}(q_{1},q_{2})$ be the unnormalized $M$-mixture of positive densities $q_{1}$ and $q_{2}$, and set $Z_{\alpha}^{M}=\int{\tilde{m}}_{\alpha}\,{\mathrm{d}\mu}$ be the normalization term so that we have $m_{\alpha}=\frac{{\tilde{m}}_{\alpha}}{Z_{\alpha}}$ and ${\tilde{m}}_{\alpha}=Z_{\alpha}m_{\alpha}$. When clear from context, we write $Z_{\alpha}$ instead of $Z_{\alpha}^{M}$.
We get after elementary calculus:
$\mathrm{JS}_{D_{{\mathrm{KL}}^{+}}}^{\tilde{M}_{\alpha},\beta}(q_{1},q_{2})=\mathrm{JS}_{D_{{\mathrm{KL}}}}^{{M}_{\alpha},\beta}(q_{1},q_{2})-(\beta Z(q_{1})+(1-\beta)Z(q_{2}))\log Z_{\alpha}+Z_{\alpha}-(\beta Z(q_{1})+(1-\beta)Z(q_{2})).$ (9)
Therefore the difference gap $\Delta_{M_{\alpha},\beta}(p_{1},p_{2})$ (written for short as $\Delta(p_{1},p_{2})$) between the normalized JSD and the unnormalized JSD between two normalized densities $p_{1}$ and $p_{2}$ (i.e., with $Z_{1}=Z(p_{1})=1$ and $Z_{2}=Z(p_{2})=1$) is
$\Delta(p_{1},p_{2}){:=}D_{\mathrm{JS}^{+}}^{\tilde{M}_{\alpha},\beta}(p_{1},p_{2})-D_{\mathrm{JS}}^{{M}_{\alpha},\beta}(p_{1},p_{2})=Z_{\alpha}-\log(Z_{\alpha})-1.$ (10)
Proposition 2 (Extended/normalize G-JSD Gap)
We have $D_{\mathrm{JS}^{+}}^{\tilde{M}_{\alpha},\beta}(p_{1},p_{2})=D_{\mathrm{JS}}^{{M}_{\alpha},\beta}(p_{1},p_{2})+Z_{\alpha}-\log(Z_{\alpha})-1$.
Thus $D_{\mathrm{JS}^{+}}^{\tilde{M}_{\alpha},\beta}(p_{1},p_{2})\geq D_{\mathrm{JS}}^{{M}_{\alpha},\beta}(p_{1},p_{2})$ when $\Delta(p_{1},p_{2})\geq 0$ and $D_{\mathrm{JS}^{+}}^{\tilde{M}_{\alpha},\beta}(p_{1},p_{2})\leq D_{\mathrm{JS}}^{{M}_{\alpha},\beta}(p_{1},p_{2})$ when $\Delta(p_{1},p_{2})\leq 0$.

### v3

@@ 1.2 Jensen–Shannon symmetrization of dissimilarities with generalized mixtures
Besides information theory [12], the JSD also plays an important role in machine learning [33, 20, 52]. However, one drawback that refrains its use in practice is that the JSD between two Gaussian distributions (normal distributions) is not known in closed-form since no analytic formula is known for the differential entropy of a two-component Gaussian mixture [34], and thus the JSD needs to be numerically approximated in practice by various methods.
To circumvent this problem, the geometric G-JSD was defined in [37] as follows:
Definition 2 (G-JSD [37])
$\mathrm{JS}_{G}(p_{1},p_{2}){:=}\frac{1}{2}\,\left({\mathrm{KL}}(p_{1},(p_{1}p_{2})_{G})+{\mathrm{KL}}(p_{2},(p_{1}p_{2})_{G})\right),$
where $(p_{1}p_{2})_{G}(x)=\frac{\sqrt{p_{1}(x)\,p_{2}(x)}}{\int\sqrt{p_{1}(x)\,p_{2}(x)}\,{\mathrm{d}\mu}}$ is the (normalized) geometric mixture of $p_{1}$ and $p_{2}$.
We have $\mathrm{JS}_{G}(p_{1},p_{2})={\mathrm{KL}}^{\mathrm{JS}}_{G}(p_{1},p_{2})$. Since by default the $M$- mixture JS-symmetrization of dissimilarities $D$ are done on the right argument (i.e., $D^{\mathrm{JS}}_{M}$), we may also consider a dual JS-symmetrization by setting the $M$-mixtures on the left argument. We denote this left mixture JS-symmetrization by $D^{\mathrm{JS}^{*}}_{M}$. We have $D^{\mathrm{JS}^{*}}_{M}(p_{1},p_{2})=({D^{*}})^{\mathrm{JS}}_{M}(p_{1},p_{2})$, i.e., the left-sided JS-symmetrization of $D$ amounts to a right-sided JS-symmetrization of the dual dissimilarity $D^{*}(p_{1},p_{2}){:=}D(p_{2},p_{1})$.
Thus a left-sided G-JSD divergence $\mathrm{JS}_{G}^{*}$ was also defined in  [37]:
Definition 3
The left-sided geometric Jensen–Shannon divergence (G-JSD) between two probability densities $p_{1}$ and $p_{2}$ is defined by
$\mathrm{JS}_{G}^{*}(p_{1},p_{2})$ ${:=}$ $\frac{1}{2}\,\left({\mathrm{KL}}((p_{1}p_{2})_{G},p_{1})+{\mathrm{KL}}((p_{1}p_{2})_{G},p_{2})\right),$
$=$ $\frac{1}{2}\,\left({\mathrm{KL}}^{*}(p_{1},(p_{1}p_{2})_{G})+{\mathrm{KL}}^{*}(p_{2},(p_{1}p_{2})_{G})\right),$
where $(p_{1}p_{2})_{G}(x)=\frac{\sqrt{p_{1}(x)\,p_{2}(x)}}{\int\sqrt{p_{1}(x)\,p_{2}(x)}\,{\mathrm{d}\mu}}$ is the (normalized) geometric mixture of $p_{1}$ and $p_{2}$.
To contrast with the numerical approximation limitation of the JSD between Gaussians, one advantage of the geometric Jensen–Shannon divergence (G-JSD) is that it admits a closed-form expression between Gaussian distributions [37]. However, the G-JSD is not anymore bounded. The G-JSD formula between Gaussian distributions has been used in several scenarii. See [16, 15, 31, 35, 48, 56, 50, 54, 23]) for a few use cases.
Proposition 3
$\mathrm{JS}_{G}(p_{1},p_{2})=\frac{1}{4}\,J(p_{1},p_{2})-B(p_{1},p_{2}),$
where $J(p_{1},p_{2}){:=}\int(p_{1}-p_{2})\log\frac{p_{1}}{p_{2}}\,{\mathrm{d}\mu}$ is Jeffreys’ divergence [25] and
$B(p_{1},p_{2})=-\log\int\sqrt{p_{1}p_{2}}\,{\mathrm{d}\mu}=-\log Z_{G}(p_{1},p_{2}),$
is the Bhattacharrya distance.
Proof:
We have:
$\mathrm{JS}_{G}(p_{1},p_{2})$ $:=$ $\frac{1}{2}\left({\mathrm{KL}}(p_{1},(p_{1}p_{2})_{G})+{\mathrm{KL}}(p_{2},(p_{1}p_{2})_{G})\right),$
$=$ $\frac{1}{2}\left(\int\left(p_{1}(x)\log\frac{p_{1}(x)\,Z_{G}(p_{1},p_{2})}{\sqrt{p_{1}(x)\,p_{2}(x)}}+p_{2}(x)\log\frac{p_{2}(x)\,Z_{G}(p_{1},p_{2})}{\sqrt{p_{1}(x)\,p_{2}(x)}}\right){\mathrm{d}\mu}(x)\right),$
$=$ $\frac{1}{2}\left(\int\left(p_{1}(x)+p_{2}(x)\right)\log Z_{G}(p_{1},p_{2})\,{\mathrm{d}\mu}(x)+\frac{1}{2}{\mathrm{KL}}(p_{1},p_{2})+\frac{1}{2}{\mathrm{KL}}(p_{2},p_{1})\right),$
$=$ $\log Z_{G}(p_{1},p_{2})+\frac{1}{4}J(p_{1},p_{2}),$
$=$ $\frac{1}{4}J(p_{1},p_{2})-B(p_{1},p_{2}).$
$\square$
Corollary 1 (G-JSD upper bound)
We have the upper bound $\mathrm{JS}_{G}(p,q)\leq\frac{1}{4}\,J(p,q)$.
Proof:
Since $B(p_{1},p_{2})\geq 0$ and $\mathrm{JS}_{G}(p_{1},p_{2})=\frac{1}{4}\,J(p_{1},p_{2})-B(p_{1},p_{2})$, we have $\mathrm{JS}_{G}(p,q)\leq\frac{1}{4}\,J(p,q)$. $\square$
@@
Although the KLD and JSD are separable divergences (i.e., $f$-divergences expressed as integrals of scalar divergences), the $M$-JSD divergence is in general not separable because it requires to normalize mixtures inside the log terms. Notice that the Bhattacharyya distance is similarly not a separable divergence but the Bhattacharyya similarity coefficient $\mathrm{BC}(p_{1},p_{2})=\exp(-B(p_{1},p_{2}))=\int\sqrt{p_{1}\,p_{2}}\,{\mathrm{d}\mu}$ is a separable “$f$-divergence”/$f$-coefficient for $f_{\mathrm{BC}}(u)=\sqrt{u}$ (here, a concave generator): $\mathrm{BC}(p_{1},p_{2})=I_{f_{\mathrm{BC}}}(p_{1},p_{2})$. Notice that $f_{\mathrm{BC}}(1)=1$, and because of the concavity of $f_{\mathrm{BC}}$, we have $I_{f_{\mathrm{BC}}}(p_{1},p_{2})\leq f_{\mathrm{BC}}(1)=1$ (hence, the term $f$-coefficient to reflect the notion of similarity measure).
1.3 Paper outline
The paper is organized as follows: We first give an alternative definition of the M-JSD in §2 (Definition 4) which extends to positive measures and do not require normalization of geometric mixtures. We call this new divergence the extended M-JSD, and we compare the two types of geometric JSDs when dealing with probability measures. In §4, we show that these normalized/extended $M$-JSD divergences can be interpreted as regularizations of the Jensen–Shannon divergence, and exhibit several bounds. We discuss Monte Carlo stochastic approximations and approximations using $\gamma$-divergences [19] in §5. For the case of geometric mixtures, although the G-JSD is not a $f$-divergence, we show that the extended G-JSD is a $f$-divergence (Proposition 5), and we express both the G-JSD and the extended G-JSD using both the Jeffreys divergence and the Bhattacharyya divergence or coefficient. We report related closed-form formula for the G-JSD and extended G-JSD between two Gaussian distributions in Section 3. Finally, we summarize the main results in the concluding section §6.
A list of notations is provided in Appendix A.
@@ 2.1 Definition and properties
  • First, we replace the KLD by the extended KLD between positive densities $q_{1}\in M_{\mu}^{+}$ and $q_{2}\in M_{\mu}^{+}$ instead of normalized densities:
    ${{\mathrm{KL}}^{+}}(q_{1},q_{2}){:=}\int\left(q_{1}\log\frac{q_{1}}{q_{2}}+q_{2}-q_{1}\right)\,{\mathrm{d}\mu},$ (10)
    (with ${\mathrm{KL}}^{+}(p_{1},p_{2})={\mathrm{KL}}(p_{1},p_{2})$) and,
  • Second, we consider unnormalized $M$-mixture densities:
    ${(q_{1}q_{2})}_{{\tilde{M}}_{\alpha}}(x){:=}{M_{\alpha}(q_{1}(x),q_{2}(x))},$
    where we use the $\tilde{M}$ tilde notation to indicate that the $M$-mixture is not normalized, instead of normalized densities ${(q_{1}q_{2})}_{M_{\alpha}}(x)$.
Consider the KLD formula between a normalized density $p_{1}$ and an unnormalized density $q_{2}=\lambda p_{1}$ for some $\lambda>1$. We have ${\mathrm{KL}}(p_{1},q_{2})=\int p_{1}\log\frac{p_{1}}{q_{2}}\,{\mathrm{d}\mu}=\int p_{1}\log\frac{1}{\lambda}\,{\mathrm{d}\mu}=-\log\lambda<0$. That is the KLD can be negative between non-normalized densities. However, the extended KLD is always guaranteed to be positive for $p_{1}>0$ and $q_{2}>0$ since it can be written as a pointwise scalar Bregman divergence integral for the negative Shannon entropy generator [28]:
${\mathrm{KL}}^{+}(p_{1},q_{2})$ $=$ $\int(p_{1}\log\frac{p_{1}}{q_{2}}+q_{2}-p_{1})\,{\mathrm{d}\mu},$
$=$ $\int B_{F}(p_{1}(x),q_{2}(x))\,{\mathrm{d}\mu}\geq 0,$
where $F(y)=y\log y-y$ is the extended Shannon negative entropy function: $B_{F}(a,b)=a\log\frac{a}{b}+b-a\geq 0$ with equality if and only if $a=b$.
The extended KLD is an extended $f$-divergence [44]: ${{\mathrm{KL}}^{+}}(q_{1},q_{2})=I_{f_{{\mathrm{KL}}^{+}}}^{+}(q_{1},q_{2})$ for $f_{{\mathrm{KL}}^{+}}(u)=-\log(u)+u-1$, where $I_{f}^{+}(q_{1},q_{2})$ denotes the $f$ divergence extended to positive densities $q_{1}$ and $q_{2}$:
$I_{f}^{+}(q_{1},q_{2})=\int q_{1}\,f\left(\frac{q_{2}}{q_{1}}\right)\,{\mathrm{d}\mu}.$
@@
As a side remark, it is preferable in practice to estimate the KLD between $p_{1}$ and $p_{2}$ by Monte Carlo methods using Eq. 10 instead of Eq. 1 in order to guarantee the non-negativeness of the KLD (Gibbs’ inequality). Indeed, the sampling of $s$ samples $x_{1},\ldots,x_{s}$, defines two unnormalized distributions $q_{1}(x)=\frac{1}{s}\sum_{i=1}^{s}p_{1}(x)\delta_{x_{i}}(x)$ and $q_{2}(x)=\frac{1}{s}\sum_{i=1}^{s}p_{2}(x)\delta_{x_{i}}(x)$ where
$\delta_{x_{i}}(x)=\left\{\begin{array}[]{ll}1,&\mbox{if $x=x_{i}$}\cr 0,&\mbox{otherwise}\end{array}\right.$.
@@
where $Z(q){:=}\int q\,{\mathrm{d}\mu}$ is the normalization factor of the positive density $q$. The divergence $\tilde{D}$} is said projective because we have for all $\lambda_{1}>0,\lambda_{2}>0$, the property that ${\tilde{D}}(\lambda_{1}q_{1},\lambda_{2}q_{2})={\tilde{D}}(q_{1},q_{2})=D^{+}(p_{1},p_{2})$ where $p_{i}=\frac{q_{i}}{Z(q_{i})}$ are the normalized densities. The projective Kullback–Leibler divergence $\widetilde{{\mathrm{KL}}}$ is thus another projective extension of the KLD to non-normalized densities which coincide with the KLD for probability densities. But the projective KLD is different from the extended KLD of Eq. 10, and furthermore we have $\widetilde{{\mathrm{KL}}}(q_{1},q_{2})=0$ if and only if $q_{1}=\lambda\,q_{2}$ $\mu$-almost everywhere for some $\lambda>0$.
@@
$D^{\mathrm{JS}^{+}}_{{\tilde{M}}_{\alpha},\beta}(q_{1},q_{2}){:=}\beta\,D^{+}\left(q_{1},{(q_{1}q_{2})}_{{\tilde{M}}_{\alpha}}\right)+(1-\beta)\,D^{+}\left(q_{1},{(q_{1}q_{2})}_{{\tilde{M}}_{\alpha}}\right),$ (11)
When $\beta=\frac{1}{2}$, we write for short $D^{\mathrm{JS}^{+}}_{{\tilde{M}}_{\alpha}}(q_{1},q_{2})$, and furthermore when $\alpha=\frac{1}{2}$, we simplify the notation to $D^{\mathrm{JS}^{+}}_{{\tilde{M}}}(q_{1},q_{2})$.
When $D^{+}={\mathrm{KL}}^{+}$, we obtain the extended geometric Jensen-Shannon divergence, $\mathrm{JS}_{{\tilde{G}}}^{+}(q_{1},q_{2})={\mathrm{KL}}^{\mathrm{JS}^{+}}_{{\tilde{G}}}(q_{1},q_{2})$:
Definition 5 (Extended G-JSD)
The extended geometric Jensen–Shannon divergence between two positive densities $q_{1}$ and $q_{2}$ is
@@
The extended G-JSD between two normalized densities $p_{1}$ and $p_{2}$ is thus
$\mathrm{JS}_{{\tilde{G}}}^{+}(p_{1},p_{2})$ $=$ $\frac{1}{2}\left(\int\left(p_{1}\log\frac{p_{1}}{\sqrt{p_{1}\,p_{2}}}+p_{2}\log\frac{p_{2}}{\sqrt{p_{1}\,p_{2}}}\right)\,{\mathrm{d}\mu}+\int\sqrt{p_{1}\,p_{2}}\,{\mathrm{d}\mu})\right)-1,$ (13)
$=$ $\frac{1}{2}\left(\int\left(p_{1}\log\sqrt{\frac{p_{1}}{p_{2}}}+p_{2}\log\sqrt{\frac{p_{2}}{{p_{1}}}}\right)\,{\mathrm{d}\mu}+Z_{G}(p_{1},p_{2})\right)-1,$ (14)
with $Z_{G}(p_{1},p_{2})=\exp(-B(p_{1},p_{2}))$.
Thus we get the following propositions:
Proposition 4
The extended geometric Jensen–Shannon divergence (G-JSD) can be expressed as follows:
$\mathrm{JS}_{{\tilde{G}}}^{+}(p_{1},p_{2})=\frac{1}{4}\,J(p_{1},p_{2})+\exp(-B(p_{1},p_{2}))-1.$
Proof:
We have
[…]
$=$ $\frac{1}{2}\,\left(\int\left(p_{1}\log\sqrt{\frac{p_{1}}{p_{2}}}+p_{2}\log\sqrt{\frac{p_{2}}{p_{1}}}+2\sqrt{p_{1}\,p_{2}}-(p_{1}+p_{2})\right)\,{\mathrm{d}\mu}\right),$
$=$ $\int\frac{1}{4}\,(p_{1}-p_{2})\log\frac{p_{1}}{p_{2}}\,{\mathrm{d}\mu}+\int\sqrt{p_{1}\,p_{2}}\,{\mathrm{d}\mu}-1,$
$=$ $\frac{1}{4}\,J(p_{1},p_{2})+\exp(-B(p_{1},p_{2}))-1.$
$\square$
Thus we can express the gap between $\mathrm{JS}_{{\tilde{G}}}^{+}(p_{1},p_{2})$ and $\mathrm{JS}_{G}(p_{1},p_{2})$:
$\Delta_{G}(p_{1},p_{2})=\mathrm{JS}_{{\tilde{G}}}^{+}(p_{1},p_{2})-\mathrm{JS}_{G}(p_{1},p_{2})=\exp(-B(p_{1},p_{2}))+B(p_{1},p_{2})-1.$
Since $Z_{G}(p_{1},p_{2})=\exp(-B(p_{1},p_{2}))$, we have:
$\Delta_{G}(p_{1},p_{2})=Z_{G}(p_{1},p_{2})-\log Z_{G}(p_{1},p_{2})-1.$
Proposition 5
The extended G-JSD is a $f$-divergence for the generator
$f_{{\tilde{G}}}(u)=\frac{1}{4}\left(u-1\right)\log u+\sqrt{u}-1.$
That is, we have $\mathrm{JS}_{{\tilde{G}}}^{+}(p_{1},p_{2})=I_{f_{{\tilde{G}}}}(p_{1},p_{2})$.
Proof:
We proved that $\mathrm{JS}_{{\tilde{G}}}^{+}(p_{1},p_{2})=\frac{1}{4}\,J(p_{1},p_{2})+\mathrm{BC}(p_{1},p_{2})-1$. The Jeffreys divergence is a $f$-divergence for the generator $f_{J}(u)=(u-1)\log u$, and the Bhattacharyya coefficient is a $f$-coefficient for $f_{\mathrm{BC}}(u)=\sqrt{u}$ (a “$f$-divergence” for a concave generator). Thus we have
$f_{{\tilde{G}}}(u)=\frac{1}{4}\left(u-1\right)\log u+\sqrt{u}-1,$
such that $\mathrm{JS}_{{\tilde{G}}}^{+}(p_{1},p_{2})=I_{f_{{\tilde{G}}}}(p_{1},p_{2})$. We check that $f_{{\tilde{G}}}(u)$ is convex since $f_{{\tilde{G}}}^{\prime\prime}(u)=\frac{\sqrt{u}(u+1)-u}{4u^{\frac{5}{2}}}$ (and by a change of variable $t=\sqrt{u}$, the numerator $t(t^{2}-t+1)$ is shown positive since the discriminant of $t^{2}-t+1$ is negative), and we have $f_{{\tilde{G}}}(1)=0$. Thus the extended G-JSD is a proper $f$-divergence. $\square$
It follows that the extended G-JSD satisfies the information monotonicity of invariant divergences in information geometry [3].
Remark 7
More generally, let us define the extended $(\alpha,\beta)$-GJSD for $\alpha\in(0,1),\beta\in(0,1)$ as
$\mathrm{JS}_{G_{\alpha},\beta}(p_{1},p_{2})=\int\left(\beta\,p_{1}\log\frac{p_{1}}{p_{1}^{\alpha}\,p_{2}^{1-\alpha}}+(1-\beta)\,p_{2}\log\frac{p_{2}}{p_{1}^{\alpha}\,p_{2}^{1-\alpha}}+p_{1}^{\alpha}\,p_{2}^{1-\alpha}-(\beta p_{1}^{\alpha}+(1-\beta)p_{2}^{1-\alpha})\right)\,{\mathrm{d}\mu}.$
Then we get the following identity:
$\mathrm{JS}_{G_{\alpha},\beta}(p_{1},p_{2})$ $=$ $\beta(1-\alpha){\mathrm{KL}}(p_{1},p_{2})+(1-\beta)\alpha{\mathrm{KL}}(p_{2},p_{1})+\mathrm{BC}_{\alpha}(p_{1},p_{2})-1.$
Furthermore, divergence $\mathrm{JS}_{G_{\alpha},\beta}$ is expressed using the $f$-divergence formula for the following generator:
$f_{\alpha,\beta}(u)=-(1-\alpha)\beta\log(u)+\alpha(1-\beta)u\log(u)+u^{1-\alpha}-(\beta+(1-\beta)u).$
Let $\alpha=\beta$. Then we have
$f_{\alpha,\alpha}^{\prime\prime}(u)=\frac{\alpha(1-\alpha)}{u^{2+\alpha}}\,(u^{\alpha}(u+1)+u)>0,\forall u>0,\forall\alpha\in(0,1).$
Hence $\mathrm{JS}_{G_{\alpha},\beta}(p_{1},p_{2})=I_{f_{\alpha,\alpha}}(p_{1},p_{2})\geq 0$, i.e., the extended $(\alpha,\alpha)$-GJSD is a $f$-divergence since $f_{\alpha,\alpha}(u)$ is strictly convex and we have $f_{\alpha,\alpha}(1)=0$.
@@
${{\mathrm{KL}}^{+}}(q_{1},q_{2}){:=}{\mathrm{KL}}(q_{1},q_{2})+\int\left(q_{2}-q_{1}\right)\,{\mathrm{d}\mu},$
although $q_{1}$ and $q_{2}$ may not need to be normalized in the $\mathrm{KL}$} term (which can then yield a potentially negative value). Letting $Z(q_{i}){:=}\int q_{i}\,{\mathrm{d}\mu}$ be the total mass of positive density $q_{i}$, we have
${{\mathrm{KL}}^{+}}(q_{1},q_{2})={\mathrm{KL}}(q_{1},q_{2})+Z(q_{2})-Z(q_{1}).$ (15)
Let ${\tilde{m}}_{\alpha}=M_{\alpha}(q_{1},q_{2})$ be the unnormalized $M$-mixture of positive densities $q_{1}$ and $q_{2}$, and set $Z_{M_{\alpha}}=\int{\tilde{m}}_{\alpha}\,{\mathrm{d}\mu}$ be the normalization term so that we have $m_{\alpha}=\frac{{\tilde{m}}_{\alpha}}{Z_{M_{\alpha}}}$ and ${\tilde{m}}_{\alpha}=Z_{M_{\alpha}}\,m_{\alpha}$. When clear from context, we write $Z_{\alpha}$ instead of $Z_{M_{\alpha}}$.
We get after elementary calculus the following identity:
$\mathrm{JS}^{+}_{\tilde{M}_{\alpha},\beta}(q_{1},q_{2})=\mathrm{JS}_{{{M}_{\alpha},\beta}}(q_{1},q_{2})-(\beta Z(q_{1})+(1-\beta)Z(q_{2}))\log Z_{\alpha}+Z_{\alpha}-(\beta Z(q_{1})+(1-\beta)Z(q_{2})).$ (16)
Therefore the difference gap $\Delta_{M_{\alpha},\beta}(p_{1},p_{2})$ (written for short as $\Delta(p_{1},p_{2})$) between the normalized JSD and the unnormalized M-JSD between two normalized densities $p_{1}$ and $p_{2}$ (i.e., with $Z_{1}=Z(p_{1})=1$ and $Z_{2}=Z(p_{2})=1$) is
$\Delta(p_{1},p_{2}){:=}{\mathrm{JS}^{+}}_{\tilde{M}_{\alpha},\beta}(p_{1},p_{2})-\mathrm{JS}^{{M}_{\alpha},\beta}(p_{1},p_{2})=Z_{\alpha}-\log(Z_{\alpha})-1.$ (17)
Proposition 6 (Extended/normalized M-JSD Gap)
The following identity holds:
${\mathrm{JS}^{+}}_{\tilde{M}_{\alpha},\beta}(p_{1},p_{2})=\mathrm{JS}_{{M}_{\alpha},\beta}(p_{1},p_{2})+Z_{\alpha}-\log(Z_{\alpha})-1.$
Thus ${\mathrm{JS}^{+}}_{\tilde{M}_{\alpha},\beta}(p_{1},p_{2})\geq\mathrm{JS}_{{M}_{\alpha},\beta}(p_{1},p_{2})$ when $\Delta(p_{1},p_{2})\geq 0$ and ${\mathrm{JS}^{+}}_{\tilde{M}_{\alpha},\beta}(p_{1},p_{2})\leq\mathrm{JS}_{{M}_{\alpha},\beta}(p_{1},p_{2})$ when $\Delta(p_{1},p_{2})\leq 0$.